# ]organization=LASIGE, Faculdade de Ciências, Universidade de Lisboa , country=Portugal

### 26 Jan 2024 v1

@@ Abstract
Knowledge sharing about emerging threats is crucial in the rapidly advancing field of cybersecurity and forms the foundation of Cyber Threat Intelligence. In this context, Large Language Models are becoming increasingly significant in the field of cybersecurity, presenting a wide range of opportunities. This study explores the capability of chatbots such as ChatGPT, GPT4all, Dolly, Stanford Alpaca, Alpaca-LoRA, and Falcon to identify cybersecurity-related text within Open Source Intelligence. We assess the capabilities of existing chatbot models for Natural Language Processing tasks. We consider binary classification and Named Entity Recognition as tasks. This study analyzes well-established data collected from Twitter, derived from previous research efforts. Regarding cybersecurity binary classification, Chatbot GPT-4 as a commercial model achieved an acceptable F1-score of 0.94, and the open-source GPT4all model achieved an F1-score of 0.90. However, concerning cybersecurity entity recognition, chatbot models have limitations and are less effective. This study demonstrates the capability of these chatbots only for specific tasks, such as cybersecurity binary classification, while highlighting the need for further refinement in other tasks, such as Named Entity Recognition tasks.
@@ 1 Introduction
Cybersecurity, a continuously evolving domain, involves experts publicly sharing their knowledge on cyberthreats. This information serves as the primary source for Cyber Threat Intelligence (CTI) tools, and researchers have contributed to the development of methods for extracting cyberthreat intelligence from text sources. With the development of Machine Learning (ML), significant advancements have been made in the field of cybersecurity. ML-based tools have deepened our understanding of cyberthreats, thereby enabling the creation of innovative and responsive solutions. A notable development in this domain is the emergence of Large Language Models (LLM), representing a substantial breakthrough in Natural Language Processing (NLP). This development has empowered chatbots to extract meaningful information about cybersecurity threats. Such advancements are exemplified by Franco et al. [1] work on SecBot, which focuses on the use of chatbots to enhance security. Arora et al. [2] is another research on developing chatbots to assess threats through sentiment analysis on social media. Moreover, Microsoft has attempted to introduce a chatbot for cybersecurity experts [3]. These efforts highlight the practical applications of chatbots in addressing cybersecurity challenges.
In today’s cybersecurity landscape, the timely detection and response to emerging threats are crucial. To address this vital requirement, we focus our research on exploring the capabilities of LLM-based chatbots and their potential to enhance cyberthreat awareness while streamlining detection processes. Specifically, we investigate the integration of LLMs into binary classification and Named Entity Recognition (NER) tasks within the scope of CTI. This integration enhances the field’s capabilities, enabling organizations to strengthen their defenses and proactively mitigate potential risks. Our work brings together the fields of LLM, NLP, and cybersecurity, resulting in a fascinating intersection where we explore the functionalities of chatbots in the security field. We evaluate the performance of chatbots as binary classifiers and in NER tasks, particularly in the context of Open Source Intelligence (OSINT) data collected from Twitter. More specifically, we use Alves et al. Twitter dataset [4] manually annotated, as a reliable source of cybersecurity-related tweets. Later, Dionisio et al. [5] annotated this dataset for the NER task. This provides an accurate testing environment that challenges chatbots in real-world situations. The LLM-based chatbots in this study include ChatGPT [6], GPT4all [7], Dolly [8], Stanford Alpaca [9], Alpaca-LoRA [10], Falcon [11], and Vicuna [12].
@@
    We present a state-of-the-art survey on chatbot models.
@@
    We examine whether the adaptive nature of chatbots allows us to shape them precisely to our preferences for cybersecurity tasks.
@@
    The study provides a comparative analysis of the practical use and performance of chatbots in specialized cybersecurity tasks, including binary classification and NER.
The remainder of this paper is organized as follows. Section 2 provides a comprehensive background on Transformers, LLMs, and their diverse applications in NLP tasks. Section 3.1 reviews related studies in the field of LLMs and their applications in chatbots. Section 4 presents a deep exploration of LLM-based chatbots and highlights their significance and capabilities. In Section 5, we shift our focus to the methodology used to evaluate chatbots, detailing our dataset, methods, and the comparison criteria employed for evaluation. Section 6 explores strategies aimed at optimizing the utilization of these chatbots, including prompt fine-tuning and text length control. Our experimental results are presented in Section 7, along with our findings. In Section 8, we focus on the challenges and limitations of this study. Section 9 thoroughly explores, analyzes, and interprets the findings. Finally, we conclude our study in Section 10 by summarizing the key contributions and insights derived from this study.
2 Background
@@ 2.1 Transformers
The introduction of transformer architecture [13] has significantly revolutionized the field of NLP. They have emerged as a preferred architecture for various NLP tasks due to their ability to overcome the limitations of previous methods such as Recurrent Neural Networks (RNNs) [14], Convolutional Neural Networks (CNNs) [15], and Long Short-Term Memory (LSTM) networks [16]. Transformer models show a unique structure that effectively captures the extensive dependencies and contextual associations of textual data. Instead of processing inputs one at a time, transformer models handle tokens or words simultaneously. This makes it easier to model the global interactions [17] and dependencies [18]. This feature makes transformers highly effective for various tasks, including but not limited to machine translation, sentiment analysis, and text generation. Transformers have emerged as the basic foundation for cutting-edge models such as Bidirectional Encoder Representations Transformers (BERT) [19] and Generative Pre-trained Transformers (GPT) [20]. These models have demonstrated exceptional performances across diverse NLP benchmarks, leading to progress in language understanding, text generation, and question answering [21]. Although transformer models have revolutionized the scope of NLP with their unique architecture and self-attention mechanisms, they are just one part of the broader landscape of LLMs.
@@ 2.2 Large Language Models
LLMs have emerged as significant innovations that revolutionize the human language process, generation, and understanding. LLMs are trained on large text datasets, including immense volumes of language data, which enables them to perform tasks that demand contextual comprehension and generate coherent and meaningful responses [22]. Within the landscape of LLMs, two widely established styles are BERT and GPT.
GPT-style models utilize a transformer architecture that effectively captures contextual dependencies and relationships within text [22]. Consequently, they have gained popularity for various applications that require NLP capabilities. BERT-style language models utilize a transformer architecture and adopt a masked language modeling objective. This approach enables them to effectively capture the contextualized meanings of words and phrases.
2.3 Leveraging LLMs for Cybersecurity NLP Tasks
LLMs have shown outstanding effectiveness in many NLP tasks including binary classification and NER. Employing the capabilities of LLMs in ML tasks, specifically within the scope of cybersecurity, offers several possibilities [23]. Two problems that stand out are text classification, which categorizes text according to its cybersecurity content, and NER, which recognizes specific cybersecurity entities in the text, for example, if it describes a vulnerability, which products are affected.
3 Related Work
In this section, we cover related studies in three categories: OSINT for cybersecurity, chatbots, and chatbots for the cybersecurity area.
@@ 3.1 OSINT for Cybersecurity
In the field of cyber security, NLP tasks such as text classification and NER, which involve exploring various cybersecurity threats and concepts, are crucial. In a previous study, SYNAPSE [24], an OSINT processing pipeline was implemented to efficiently identify and concisely present various cybersecurity incidents to security analysts. SYNAPSE was inspired by the results of [4], which analyzed and discussed the publication of OSINT on Twitter. Unlike models based on modern Deep Learning (DL) methods, SYNAPSE utilizes K-means clustering and Support Vector Machines (SVM) [25] for its operations. To improve SYNAPSE, Dionisio et al. [26] showcased the workflow of a new tool that employs a CNN to detect cybersecurity texts gathered from Twitter. That work was further improved by a multitask DL approach [5] that could simultaneously classify and perform NER on these tweets.
3.2 Chatbots
Although traditional and DL techniques have paved the way for advancements in cybersecurity text classification and NER, a new wave of NLP tools offers even more refined capabilities. Among these, LLMs stand out for their exceptional capabilities in interpreting and generating human-like text. [22] presents the practical LLM utilized in NLP tasks. A complete guide explains how to use them from various perspectives including model selection, data consideration, and task specificity. It thoroughly investigates the details of pre-training and the significance of the training and test data. It offers insights into tasks that require a rich knowledge base, natural language comprehension, generation capabilities, and other emergent features [22]. Concurrently, LLMs, such as the Generative Pre-trained Transformer 3 (GPT-3), have gained attention for their remarkable ability to acquire a broad spectrum of knowledge during pre-training and apply it to downstream NLP tasks [27], with significant potential for chatbots [28, 12, 8, 9, 10, 11, 6]. Chatbots powered by LLMs have attracted interest as powerful tools for data annotation in NLP domains [27]. This interest arises from chatbots’ proficiency in language tasks and the critical role of data annotation in the development of NLP systems. An illustrative study [29] compared ChatGPT with human crowd workers in annotation tasks. The research emphasized that ChatGPT surpassed human workers in terms of performance, agreement among annotators, and cost-effectiveness.
However, it is essential to consider the strengths and challenges of LLM chatbots. An analysis of ChatGPT’s performance [30] showed that despite its achievements, it frequently fell behind supervised baselines in various NLP tasks. Several factors influence this, including limitations on the number of tokens, a misalignment with specific NLP tasks due to its generative nature, and challenges inherent to LLMs, such as hallucination, which involves making false positive predictions. In their effort to optimize ChatGPT, researchers have proposed solutions, including multiple prompts, task-specific fine-tuning, and strategies to counter hallucination [30]. These methods were thoroughly tested across 21 datasets, covering ten critical NLP tasks, including the NER task. The testing resulted in significant performance improvements for ChatGPT, with instances where it outperformed state-of-the-art models existing in benchmarks [30].
The study by Megahed et al. [31] reveals that ChatGPT excels in structured tasks like code translation and explaining established concepts. However, it faces challenges when handling nuanced tasks such as recognizing unfamiliar entities and generating code from scratch. Recent research findings show that the ChatGPT may encounter challenges in accurately identifying popular entities, including locations, names, and organizations. Therefore, their performance faces clear limitations when it comes to recognizing specific entities, particularly in the context of security-related text [32]. In the research [32], the authors present the results of their experiments on the performance of GPT-3.5, ChatGPT, and fine-tuned models on the multi-domain CONLL dataset for recognizing entities. According to the findings reported in this paper, ChatGPT and GPT-3.5 achieved F1-score of 53.7% and 53.5%, respectively. In addition, Xiaofei et al. [30] investigated the factors contributing to the suboptimal performance of GPT-based chatbots in NLP tasks, such as NER. They have identified several underlying causes and have proposed a set of generalized modules to mitigate these challenges in different NLP tasks. These efforts aim to enhance the capabilities of ChatGPT and extend its applicability to a broader range of NLP tasks.
@@ 3.3 Chatbots for Cybersecurity
As LLM-based chatbots have become widespread, concerns about cybersecurity threats and vulnerabilities have increased [33]. The research examines the history of chatbot development, how LLMs operate, and why it is crucial to address cyberattacks and vulnerabilities to reduce risks [33]. Moreover, ChatGPT can serve as a potential honeypot interface for cybersecurity by imitating terminal commands and providing an interface for TeamViewer, Nmap, and Ping. This dynamic environment can adapt to the actions of attackers and provide insights into their tactics, techniques, and procedures. The usefulness of ChatGPT extends the time-to-conquer or delay attacker timelines, making it valuable for organizations seeking to enhance their cybersecurity posture [34]. Additionally, the performances of the ChatGPT and GPT-3 models is evaluated for vulnerability detection in the code. Based on a real-world dataset, this evaluation focuses on binary and multi-label classification tasks related to Common Weakness Enumeration (CWE) vulnerabilities. The findings indicate that ChatGPT does not outperform the baseline classifier in binary and multi-label classification tasks for code vulnerability detection [35].
@@
Sentiment analysis plays a crucial role in the extraction of user opinions and emotions from textual data to assess threats. Recent research has been directed toward developing sustainable strategies to diminish threats, vulnerabilities, and data manipulation within chatbots, ultimately improving the scope of cybersecurity. To achieve this objective, researchers created an interactive chatbot and placed it on social media platforms such as Twitter for the specific purpose of cybersecurity [2]. This study employs a sentiment analysis strategy by deploying chatbots on Twitter and subsequently analyzing Twitter data to anticipate forthcoming threats and cyberattacks.

### 13 Mar 2024 v2

@@ Abstract
Knowledge sharing about emerging threats is crucial in the rapidly advancing field of cybersecurity and forms the foundation of Cyber Threat Intelligence (CTI). In this context, Large Language Models are becoming increasingly significant in the field of cybersecurity, presenting a wide range of opportunities. This study surveys the performance of ChatGPT, GPT4all, Dolly, Stanford Alpaca, Alpaca-LoRA, Falcon, and Vicuna chatbots in binary classification and Named Entity Recognition (NER) tasks performed using Open Source INTelligence (OSINT). We utilize well-established data collected in previous research from Twitter to assess the competitiveness of these chatbots when compared to specialized models trained for those tasks. In binary classification experiments, Chatbot GPT-4 as a commercial model achieved an acceptable $\text{F}_{1}$ score of 0.94, and the open-source GPT4all model achieved an $\text{F}_{1}$ score of 0.90. However, concerning cybersecurity entity recognition, all evaluated chatbots have limitations and are less effective. This study demonstrates the capability of chatbots for OSINT binary classification and shows that they require further improvement in NER to effectively replace specially trained models. Our results shed light on the limitations of the LLM chatbots when compared to specialized models, and can help researchers improve chatbots technology with the objective to reduce the required effort to integrate machine learning in OSINT-based CTI tools.
@@ 1 Introduction
Cybersecurity, a continuously evolving domain, involves experts publicly sharing their knowledge on cyber threats. This information is the primary source for Cyber Threat Intelligence (CTI) tools, and researchers have contributed to developing methods for extracting cyber threat intelligence from text sources (e.g., [1, 2]). With the development of Machine Learning (ML), significant advancements have been made in the field of cybersecurity. ML-based tools have deepened our understanding of cyber threats, enabling innovative and responsive solutions. A notable development in this domain is the emergence of Large Language Models (LLM), representing a substantial breakthrough in Natural Language Processing (NLP). This development has empowered chatbots to extract meaningful information about cybersecurity threats. Such advancements are exemplified by academic works such as SecBot [3] and Arora et al. [4], which provide proof-of-concept chatbots to support cybersecurity planning and management, and recent high-profile products such as Microsoft Security Copilot [5]. These efforts (among others, described in Section 2) highlight the practical use of chatbots to address cybersecurity and CTI challenges.
In today’s cybersecurity landscape, the timely detection and response to emerging threats are crucial. To address this vital requirement, we focus on researching the potential of LLM-based chatbots to improve cyber threat awareness and streamline the detection processes. Specifically, we investigate using pre-trained LLM chabots for binary classification and Named Entity Recognition (NER) tasks within the scope of generating indicators of compromise for CTI. This integration enhances the field’s capabilities, enabling organizations to strengthen their defences and proactively mitigate potential risks. The important question remains of whether LLM-based chatbots can achieve the level of performance [6] of specialized models [7], which are known to achieve excellent results using a wide variety of techniques for binary classification [8], and NER [9] tasks.
This paper presents an empirical study that follows a comparative approach to answer the following research question: are LLM chatbots competitive with state-of-the-art specialized models for detecting OSINT CTI and extracting pertinent information? To produce the empirical results and evaluate the performance of chatbots, we use a publicly-available annotated dataset [10, 11] collected from Twitter, a known reliable OSINT CTI source [12]. This dataset provides an accurate testing environment for challenging chatbots in a real-world CTI scenario.
The LLM-based chatbots in this study include ChatGPT [13], GPT4all [14], Dolly [15], Stanford Alpaca [16], Alpaca-LoRA [17], Falcon [18], and Vicuna [19]. These systems represent the two prominent types of chatbots widely used nowadays: paid chatbots available as a service through APIs, and Open-source chatbot models that are built to run on local GPU servers. When planning the experiments, we picked all the available variants of GPT-like models since they are standard in the field [20].
Since utilizing LLM chatbots raises additional technical questions, we also study the impact of different chatbot utilization methods on the empirical results obtained to answer the main research question. These questions concern the capability of chatbots to provide clear yes-or-no answers, and the cost (in terms of processing time) required for these chatbots to perform the NLP tasks.
@@
    We present a state-of-the-art survey on chatbot models and their application to cybersecurity.
@@
    We investigate the extent to which the inherent flexibility of LLM-based chatbots can be tailored to meet the specific requirements of OSINT-based CTI applications.
@@
    The study provides a comparative analysis of the practical use and performance of chatbots in specialized CTI tasks, including binary text classification and NER.
The remainder of this paper is organized as follows. Section 2 provides some background on LLMs and chatbots, and reviews related works on LLM chatbots for cybersecurity, and the evaluation of chatbots in NLP tasks. Section 3 presents a deep exploration of LLM-based chatbots and highlights their significance and capabilities. In Section 4, we shift our focus to the methodology used to evaluate chatbots, detailing our dataset, methods, and the comparison criteria employed for evaluation. Section 5 explores strategies aimed at optimizing the utilization of these chatbots, including prompt fine-tuning and text length control. Section 6 presents our main experimental results and their discussion. Finally, we conclude our study in Section 7 by summarizing the key contributions and insights derived from this study.
2 Background and Related work
In this section, we briefly present the required background for this paper. We start by analyzing transformer models, the foundational elements facilitating recent progress in NLP. Next, we discuss LLMs by examining their development, abilities, and significant influence on various specific areas. Building on this foundation, we review research on how LLMs can effectively be utilized for cybersecurity-related NLP, exploring in detail the role of OSINT in cybersecurity. We end the section by examining chatbots’ evolution and NLP capabilities to address cybersecurity concerns and the literature gap we aim to contribute with this paper.
@@ 2.1 Transformers
The introduction of transformers [21] revolutionized the field of NLP, as they became the preferred architecture for various NLP tasks due to their ability to effectively capture the extensive dependencies and contextual associations of textual data [22, 23]. This ability enables transformers to overcome the limitations of previous methods such as Recurrent Neural Networks (RNNs) [24], Convolutional Neural Networks (CNNs) [25], and Long Short-Term Memory (LSTM) [26]. Instead of processing inputs one at a time, transformer models handle tokens or words simultaneously. This makes it easier to model the global interactions [27, 28] and dependencies [23]. This feature makes transformers highly effective for various tasks, including but not limited to machine translation, sentiment analysis, and text generation. Transformers have emerged as the basic foundation for models such as Bidirectional Encoder Representations Transformers (BERT) [29] and Generative Pre-trained Transformers (GPT) [30]. These models have demonstrated exceptional performances across diverse NLP benchmarks, leading to progress in language understanding, text generation, and question answering [31].
@@ 2.2 Large Language Models
LLMs have emerged as significant innovations that revolutionize the human language process, generation, and understanding. These models are trained on large text datasets, including immense volumes of language data, which enables them to perform tasks that demand contextual comprehension and generate coherent and meaningful responses [32]. In terms of applications, the latest generation of language models has been applied across diverse fields. For instance, the utilization of LLMs within mathematics, physics, and chemistry problem-solving has been evaluated by Arora et al. [33]. Agrawal evaluated the ability of LLMs for human-like reasoning tasks [34]. The findings of the study indicate that LLMs have strong abilities in analogical and moral reasoning but face challenges in spatial reasoning tests. Chatbots powered by LLMs have attracted interest as powerful tools for data annotation in NLP domains [35]. This interest arises from chatbots’ proficiency in language tasks and the critical role of data annotation in developing NLP systems. An illustrative study [36] compared ChatGPT with human crowd workers in annotation tasks. The research emphasized that ChatGPT surpassed human workers in terms of performance, agreement among annotators, and cost-effectiveness.
There are two widely used types of LLMs: GPT and BERT. GPT-style models utilize a regressive transformer architecture to capture contextual dependencies and relationships within text [32], while BERT-style language models utilize a bidirectional transformer architecture and adopt a masked language modelling objective. Both approaches gained popularity for NLP capabilities because of their ability to model contextual dependencies. Besides the difference in the context they consider to make text predictions, they also differ in the number of parameters and training dataset size, with GPT-style models requiring more resources in both cases.
Yang et al. [32] provides a complete guide explaining how to use LLMs from various perspectives, including model selection, data consideration, and task specificity. It thoroughly investigates the details of pre-training and the significance of the training and test data and offers insights into tasks that require a rich knowledge base, natural language comprehension, generation capabilities, and other emergent features. In particular, LLMs such as the Generative Pre-trained Transformer 3 (GPT-3), have gained attention due to their remarkable ability to acquire a broad spectrum of knowledge during pre-training and apply it to downstream NLP tasks [35], with significant potential for chatbots [37, 19, 15, 16, 17, 18, 13].
2.3 Leveraging LLMs for cybersecurity NLP
Employing the capabilities of LLMs in ML tasks, specifically within the scope of cybersecurity, offers several possibilities [38]. Two problems that stand out are text classification, which categorizes text according to its relevance to a cybersecurity context, and NER, which recognizes specific cybersecurity entities in the text, for example, if it describes a vulnerability, which products are affected. LLMs have shown outstanding effectiveness in many NLP tasks, including binary classification and NER [31, 6].
@@ 2.4 OSINT for cybersecurity
In the CTI field, NLP tasks such as text classification and NER, which involve exploring various cybersecurity threats and concepts, are crucial. In a previous study, SYNAPSE [10], an OSINT processing pipeline was implemented to efficiently identify and concisely present various cybersecurity incidents to security analysts. SYNAPSE was motivated by the results of previous works (e.g., [12, 39]) that analyzed the completeness and timeliness of cybersecurity-related OSINT on Twitter.
SYNAPSE employed Support Vector Machines (SVM) [40] for binary classification, and a novel stream clustering approach to aggregate related tweets. To improve SYNAPSE, Dionísio et al. [11] designed a new tool that employs a CNN to detect cybersecurity-related texts gathered from Twitter and a BiLSTM network for performing NER on the detected tweets to identify the type of threat, affected products and other information. That work was further extended using a multitask Deep Learning (DL) approach [41] that could simultaneously perform classification and NER.
Although there are many ML techniques for binary classification and NER [8, 9], to the best of our knowledge, the architecture of Dionísio et al. [11, 41] is recognized [42] as the state-of-the-art for OSINT-based CTI extraction on Twitter data, reaching almost 95% for different quality metrics. This work also shows superior performance when compared with classifiers based on SVMs and Multi-Layer Perceptrons (MLPs), and several alternative NER approaches. Due to this, we consider Dionisio et al. [41] as the reference specialized model in our comparison.
@@ 2.5 Chatbots for Cybersecurity
Although traditional and DL techniques have paved the way for advancements in cybersecurity text classification and NER, a new wave of NLP tools offers even more refined capabilities. Among these tools, LLMs stand out for their exceptional capabilities in interpreting and generating human-like text.
The specific case of ChatGPT raised discussions and reviews on its potential applications in the cybersecurity field [43, 44, 45] and specific attack mitigation approaches have already been described [46]. Other works designed chatbots specifically to help cybersecurity analysis. SecBot [3] is a cybersecurity-driven conversational chatbot that extracts information from a conversation to support cybersecurity planning and management. It was developed and evaluated on a small dataset utilizing the Rasa framework [47], achieving 100% accuracy in extracting the intent of attacks and associated named entities.
2.6 Evaluation of LLM Chatbots
Given the widespread attention LLM chatbots are attracting, several works tried to evaluate the strengths and limitations of the technology. An analysis of ChatGPT’s performance [6, 48] showed that despite its achievements, it frequently fell behind supervised baselines in various NLP tasks. Several factors influence this, including limitations on the number of tokens, a misalignment with specific NLP tasks due to its generative nature, and challenges inherent to LLMs, such as hallucination, which involves making false positive predictions. In their effort to optimize ChatGPT, the authors have proposed solutions, including multiple prompts, task-specific fine-tuning, and strategies to counter hallucination. These methods were thoroughly tested across 21 datasets, covering ten critical NLP tasks, including NER. The testing of the solutions resulted in significant performance improvements for ChatGPT, with instances where it outperformed state-of-the-art models in existing benchmarks [48].
The study by Megahed et al. [49] shows that ChatGPT excels in structured tasks like code translation and explaining established concepts. However, it faces challenges when handling nuanced tasks such as recognizing unfamiliar entities and generating code from scratch.
Recent research findings show that the ChatGPT may encounter challenges and limitations in accurately identifying entities, including locations, names, and organizations. Qin et al. [50] present the results of their experiments on the performance of GPT-3.5, ChatGPT, and fine-tuned models on the multi-domain CONLL dataset for recognizing entities. According to the findings reported in this paper, ChatGPT and GPT-3.5 achieved $\text{F}_{1}$ score of 53.7% and 53.5%, respectively. Sun et al.  [48] investigated the factors contributing to the sub-optimal performance of GPT-based chatbots in NLP tasks, such as NER. They have identified several underlying causes and have proposed a set of generalized modules to mitigate these challenges in different NLP tasks. Kocoń et al. [6] examined the capabilities of ChatGPT on a diverse set of subjective analytical tasks and objective reasoning tasks, revealing that when compared to state-of-the-art models, the average quality loss is about 25% in zero-shot and few-shot settings.
@@ 2.6 Evaluation of LLM Chatbots
As LLM-based chatbots have become widespread, concerns about their own vulnerabilities and their role as tools for cyber- attacks have increased [51]. The usefulness of ChatGPT extends the time-to-conquer or delay attacker timelines, making it valuable for organizations seeking to enhance their cybersecurity posture [52]. Additionally, the performances of the ChatGPT and GPT-3 models are evaluated for vulnerability detection in code [53]. Based on a real-world dataset, this evaluation focuses on binary and multi-label classification tasks related to common weakness enumeration of vulnerabilities. The findings indicate that ChatGPT does not outperform the baseline classifier in classification tasks for code vulnerability detection [53]. More precisely, the performance of both GPT models was assessed by accuracy, precision, recall, $\text{F}_{1}$ score, and area under the curve. The highest $\text{F}_{1}$ score of 0.67 was achieved using the text-davinci-003 model for binary classification. In contrast, the $\text{F}_{1}$ score of all ChatGPT models remained below 0.53 for multilabel classification.
@@
Sentiment analysis plays a crucial role in extracting user opinions and emotions from textual data to assess threats. Recent research has been directed toward developing sustainable strategies to diminish threats, vulnerabilities, and data manipulation within chatbots, ultimately improving the scope of cybersecurity. To achieve this objective, researchers created an interactive chatbot using the Bot Libre platform [^1] https://www.botlibre.com/. and placed it on social media platforms such as Twitter for the specific purpose of cybersecurity [4]. This study employs a sentiment analysis strategy by deploying chatbots on Twitter and subsequently analyzing Twitter data to anticipate forthcoming threats and cyberattacks.
2.7 The Research Gap